OpenStudy (anonymous):

You need to prepare 100.0 mL of a pH=4.00 buffer solution using 0.100 M benzoic acid (pKa = 4.20) and 0.140 M sodium benzoate.

OpenStudy (aravindg):

We know that for a buffer solution pH = pKa + log [salt]/[acid] 4 = 4.20 + log [salt] / [acid] -0.20 = log [salt]/[acid] Solve for [salt]/[acid] This gives ratio of moles. Take amount of acid be x L and then amount of salt will be 0.1-x Now molarity*volume=number of moles Okay I have given a basic idea. Use the details to get to the result. Let me know if you are still facing difficulty.